OpenStudy (anonymous):

Evaluate the integral Int((t)(e^-t))dt from 0 to 1 (Integration by parts)

OpenStudy (unklerhaukus):

\[\int\limits_0^1\ln(t)e^{-t}\text dt\]

OpenStudy (anonymous):

No ln(t) just t

OpenStudy (anonymous):

\[\int\limits_{0}^{1}t(e ^{-t})dt\]

OpenStudy (unklerhaukus):

oh, like this? \[\int\limits_0^1 (t)e^{-t}\text dt\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

yeah ok i see now

OpenStudy (unklerhaukus):

Integration by parts\[\int\limits_a^b u\text d v=[uv]_a^b-\int\limits_a^b v\text du\]

OpenStudy (anonymous):

yes but I'm having trouble evaluating it

OpenStudy (unklerhaukus):

choose \(u=t\qquad\text du=\text dt\) \[\text dv=e^{-t}\qquad v={-e^{-t}}\]

OpenStudy (anonymous):

I did, I'm still having trouble - would you possibly be able to do the whole thing out for me?

OpenStudy (unklerhaukus):

\[\int\limits_a^b u\text d v=[uv]_a^b-\int\limits_a^b v\text du\]\[\downarrow\]\[\int\limits_0^1 te^{-t}\text dt=[-te^{-t}]_0^1+\int\limits_0^1 e^{-t}\text dt\]